At fiscal year end, December 31, 2018, Somerset Corporation had total stockholders' equity of $4,600,000. On FY 2018 year end, Somerset Corporation had Common Stock account of $1,500,000 of $10 par value common stock and Preferred Stock account of $200,000 of $100 par value.  There was no treasury stock. The preferred stock was noncumulative and had a call price of $102.  Use this information to determine the book value per share of Common Stock as of end of the FY 2018: (Round your answer to the nearest penny.)
Answer #1

No. Of common stock = 1500000/10 = 150000

Call value of preferred stock = 200000*102% = $204000

Total common stock holders equity = 4600000 - 204000

= $4396000

Book value per share of Common stock = $4396000/150000

= $29.31 per share